# AlphaOPT: Formulating Optimization Programs with Self-Improving LLM Experience Library

### v2

@@ Abstract
Optimization modeling enables critical decisions across industries but remains hard to automate: informal language must be mapped to precise mathematical formulations and executable solver code, while prior LLM approaches either rely on brittle prompting or costly retraining with limited generalization. We present AlphaOPT, a self-improving experience library that enables an LLM to learn from limited demonstrations (i.e, even answers alone without gold-standard program) and solver feedback without annotated reasoning traces or parameter updates. AlphaOPT operates a continual two-phase cycle: (i) a Library Learning phase that reflects on failed attempts, extracts solver-verified, structured insights as $\{\textit{taxonomy},\ \textit{condition},\ \textit{explanation},\ \textit{example}\}$; and (ii) a Library Evolution phase that diagnoses retrieval misalignments and refines the applicability conditions of stored insights, improving transfer across tasks. This design (1) learns efficiently from limited demonstrations without curated rationales, (2) expands continually without costly retraining by updating the library rather than model weights, and (3) makes knowledge explicit and interpretable for human inspection and intervention. Experiments show that AlphaOPT steadily improves with more data (65% $\rightarrow$ 72% from 100 to 300 training items) and surpasses the strongest baseline by 7.7% on the out-of-distribution OptiBench dataset when trained only on answers. AlphaOPT code and data are available at https://github.com/Minw913/AlphaOPT.
1 INTRODUCTION
@@
Optimization models support critical decision-making in finance, manufacturing, marketing, transportation, and logistics (AhmadiTeshnizi et al., 2023; Bertsimas & Tsitsiklis, 1997; Ramamonjison et al., 2022). Beyond improving efficiency, automating the optimization workflow lowers the barrier to operations research expertise in industry, enabling non-experts to prototype faster, iterate on formulations, and deploy solver-backed decisions at scale. Yet this process has long been challenging, as informal and often ambiguous specifications must be mapped to precise, domain-specific formulations and paired with appropriate code and solvers, creating major bottlenecks for end-to-end automation (Jiang et al., 2025).
@@
Advances in large language models (LLMs) make this vision increasingly feasible: they can parse natural language requirements (Ouyang et al., 2022), generate executable programs (Nijkamp et al., ; Jimenez et al., 2024), and orchestrate downstream tools (Qin et al., 2024). Two main lines of work have emerged. Prompt-based systems steer general LLMs with structured prompts and tool use (Xiao et al., 2023; Thind et al., 2025; AhmadiTeshnizi et al., 2024; Zhang & Luo, 2025). Fine-tuning approaches adapt models on domain corpora and benchmarks (Huang et al., 2025; Yang et al., 2024). Despite this progress, both families face limitations: prompt-based systems stop improving once they run out of fixed templates, and they are fragile to small wording changes and shifts in the domain; fine-tuned models require costly retraining and, critically, most benchmarks and datasets in the community (e.g., NLP4LP (AhmadiTeshnizi et al., 2024), MAMO (Huang et al., 2024), IndustryOR (Huang et al., 2025)) contain only programs/solutions rather than the intermediate reasoning that governs modeling choices, thereby limiting the generalizability of fine-tuning approaches. This motivates a new learning paradigm for optimization formulation: instead of relying solely on prompts or retraining, LLMs should continually improve by accumulating, refining, and reusing solver-verified modeling insights.
Figure 1: AlphaOPT identifies and retrieves relevant insights to guide problem solving. In this example, it avoids the common mistake of minimizing the sum of process times and instead introduces an auxiliary variable to correctly minimize the makespan, leading to the correct solution.
We propose AlphaOPT, a self-improving framework that builds and refines a structured library of solver-verified insights for optimization formulation, as exemplified in Figure 1. Each insight encodes a reusable modeling rule in the form of a 4-tuple (taxonomy, (applicability) condition, explanation, example), which specifies not only what to reuse but also when and why it applies. We remark that our library learning framework does not require backpropagation to update framework parameters and can be regarded as the evolutionary mechanism. AlphaOPT improves through a continual two-phase cycle. Library Learning acquires new insights from both gold programs (when available) and solver-verified answer-only supervision, organizing them into a dynamically updating hierarchical taxonomy. Library Evolution then diagnoses misalignments between tasks and insight applicability, and refines conditions using aggregate evidence, ensuring that insights remain neither too narrow nor overly general. This design yields a principled optimization view: library construction corresponds to maximizing expected task success induced by task–insight matching while regularizing size to maintain efficiency and prevent redundancy.
We conduct quantitative experiments across multiple benchmarks and baselines, as well as qualitative analyses of the learned library. The results show that, compared to prompt-based or fine-tuning approaches, AlphaOPT (1) learns efficiently from limited demonstrations (i.e., it can learn from answers without recalling formulation) without requiring annotated reasoning traces or even gold-standard programs, (2) achieves stronger out-of-distribution generalizability and more consistent continual growth than learning-based methods, and (3) makes knowledge explicit and interpretable for human inspection and involvement.
@@
Beyond these advantages, AlphaOPT also achieves state-of-the-art performance on multiple benchmarks. These results demonstrate the efficacy and potential of self-improving experience-library learning for optimization formulation, paving the way toward more challenging settings, such as efficient program formulation and large-scale optimization.
Our main contributions are threefold:
  • A library learning framework that learns from answers only. We propose the first experience-library learning framework for natural language optimization formulation tasks, formally grounded in a mathematical view. The system can learn solely from answers, without requiring gold-standard programs.
  • A reusable and interpretable experience library. We construct the first solver-verified library of structured modeling insights for LLM systems, designed to be reusable across tasks and explicitly interpretable for reliable transfer in operations research domains.
@@
  • State-of-the-art out-of-distribution generalization. AlphaOPT achieves strong generalization beyond training distributions, attaining state-of-the-art performance on LogiOR and OptiBench benchmarks.
2 Related Work
LLMs for Solving Optimization Problems. Related work can be categorized into prompt-based and learning-based approaches. Prompt-based methods guide reasoning and modeling through multi-step prompts using proprietary LLMs (AhmadiTeshnizi et al., 2024; Xiao et al., 2023). AhmadiTeshnizi et al. (2023) first introduced OptiMUS, demonstrating how LLMs can generate optimization formulations from natural language descriptions, and OptiMUS-0.3 (AhmadiTeshnizi et al., 2024) extends this line of work to large-scale instances, introducing retrieval-augmented prompting and solver-integrated verification to improve scalability.
In contrast, learning-based methods construct training datasets and apply instruction tuning to open-source LLMs. Huang et al. (2025) proposed a semi-automated pipeline to synthesize training data, which is then used to fine-tune an open-source ORLM model. LLMOPT (Jiang et al., 2025) combines both paradigms by modeling optimization with five elements and fine-tuning on expert-annotated data via multi-instruction SFT. More recently, ORThought (Yang et al., 2025a) introduced the LogiOR benchmark and an expert-guided chain-of-thought framework, providing a systematic dataset and evaluation pipeline for optimization tasks that require harder logic. In terms of multi-agent design, Xiao et al. (2023) employs a collaborative multi-expert framework to enhance reasoning, Zhang & Luo (2025) integrates sandbox-based code execution and self-repair/self-verification.
Several benchmarks now exist that cover LP, MILP, NLP, and other optimization problem types (Xiao et al., 2023; AhmadiTeshnizi et al., 2024; Huang et al., 2025; Yang et al., 2024). Yet, none of the prior work has investigated strengthening LLMs’ optimization capabilities by learning and reusing structured modeling experience.
Decision-making tasks with Library Learning. Library Learning refers to the process where reusable patterns or modules are automatically extracted from past experiences to improve future problem-solving. These experiences include concrete trajectories or demonstrations, as well as abstracted rules generalized from successful or failed attempts (Zhao et al., 2024; Mu et al., 2025; Feng et al., 2025; Wang et al., ; Zhu et al., 2023). In terms of experience improvement, Zhao et al. (2024) and Mu et al. (2025) leverage an LLM to prune the library by checking if a newly added insight duplicates or conflicts with existing insights, or merges and generalizes from those overlapping insights. Feng et al. (2025) uses check functions to ensure that LLM-translated action sequences remain within the generalization boundaries of the original experiences.
LLM-driven Evolutionary Methods. Recent LLM-driven evolutionary frameworks have achieved remarkable advances in scientific discovery, showcasing LLM’s capacity for broad generative exploration on solutions. Romera-Paredes et al. (2024) introduces FunSearch, a genetic programming driven by LLMs to search for feasible or optimal solutions of mathematical problems. AlphaEvolve (Novikov et al., 2025) extends the FunSearch system to provide the ability to perform multiobjective optimization using rich forms of natural-language context and feedback. Grayeli et al. (2024) applies LLMs to discover abstract concepts from high-performing hypotheses, combining symbolic regression with LLM-guided exploration within a concept library. ReEvo (Ye et al., 2024) frames LLMs as hyper-heuristics with a reflective evolution mechanism, enabling the generation and iterative refinement of heuristics across multiple NP-hard problems. HeurAgenix (Yang et al., 2025b) further develops this direction by evolving a pool of heuristics and dynamically selecting the most suitable one for each problem state, highlighting LLMs’ role in adaptive heuristic discovery. Besides, LLM-guided evolution has also found use in discovering heuristics for combinatorial optimization (Liu et al., ).
3 Methodology
@@
Optimization tasks arrive with diverse natural-language descriptions, yet they share recurring modeling rules that activate under identifiable conditions. We identify three major challenges in building reliable systems that leverage LLMs to formulate and solve optimization problems using existing technologies and resources. First, gold-standard programs are scarce and may contain annotation errors (Jiang et al., 2025; Yang et al., 2025a), while datasets with only answer labels remain underutilized (Huang et al., 2024; 2025; Lu et al., 2025). Second, fine-tuned models (Huang et al., 2025; Jiang et al., 2025) struggle to generalize because the crucial when-to-apply-what knowledge is weakly represented (or missing) in training data; they can mimic syntax without mastering applicability. Third, the performance of prompt-based agent systems AhmadiTeshnizi et al. (2023); Xiao et al. (2023); Yang et al. (2025a) stagnates as the number of exemplars increases: they rely on human empirical curation and lack the capacity to adapt or to continually learn from larger datasets.
We propose AlphaOPT, an experience-library learning framework that learns from both gold programs (when available) and answer-only supervision. AlphaOPT iteratively builds a structured, solver-verified repository of reusable insights with explicit applicability conditions and evolves these conditions at the population level to improve generalization while avoiding redundancy. This two-stage design is described in Section 3.1. In Section 3.2, we provide a mathematical interpretation that frames library construction as maximizing task success with a size regularizer. In Section A, we compare our method with prior works on learning from experience and self-evolving problem-solving agents.
3.1 AlphaOPT Framework
Figure 2: AlphaOPT builds and refines an experience library over multiple iterations. Each iteration consists of a library learning stage, which produces insights from attempts to solve tasks, and a library refinement stage, which adjusts the applicability of insights to avoid being too narrow or too general.
The framework incrementally learns a structured library of experiences over iterations until a stopping criterion indicates that the current model can no longer make meaningful improvements. As illustrated in Figure 2, each iteration consists of two complementary phases that form a continual cycle of acquisition and refinement. The first phase, Library Learning, extracts insights from individual tasks under either gold-program or answer-only supervision while minimizing redundancy. The second phase, Library Evolution, diagnoses misalignments between insights and tasks and refines applicability conditions to enhance generalization while reducing confusion caused by overgeneralization. The design follows three guiding principles: it is failure-driven (every error becomes a learning opportunity), locally validated (an insight must solve its source task before being admitted), and compact yet generalizable (redundant insights are merged and conditions refined to prevent uncontrolled growth that hinders retrieval and execution). The prompts for all LLM modules are provided in Appendix E.
@@ 3.1 AlphaOPT Framework
Figure 3: The overall workflow of library learning and evolution. The left panel depicts two complementary flows: library learning, which extracts new insights by generating and consolidating insights from failed optimization attempts, and library diagnosis, which analyzes interactions between failed tasks and retrieved insights to collect negative and unretrieved cases for refinement. The right panel illustrates library refinement, where the LLM refines each insight’s applicability conditions, validates them via retrieval replay, and integrates the updated insights back into the library.
3.1.1 Library Learning

### v3

# AlphaOPT: Formulating Optimization Programs with a Self-Improving LLM Experience Library

@@ Abstract.
Optimization modeling underlies critical decision-making across industries, yet remains difficult to automate: natural-language problem descriptions must be translated into precise mathematical formulations and executable solver code. Existing LLM-based approaches typically rely on brittle prompting or costly retraining, both of which offer limited generalization. Recent work suggests that large models can improve via experience reuse, but how to systematically acquire, refine, and reuse such experience in structurally constrained settings remains unclear. We present AlphaOPT, a self-improving experience library that enables LLMs to learn optimization modeling knowledge from limited supervision, including answer-only feedback without gold-standard programs, annotated reasoning traces, or parameter updates. AlphaOPT operates in a continual two-phase cycle: a Library Learning phase that extracts solver-verified, structured insights from failed attempts, and a Library Evolution phase that refines the applicability of stored insights based on aggregate evidence across tasks. This design allows the model to accumulate reusable modeling principles, improve transfer across problem instances, and maintain bounded library growth over time. Evaluated on multiple optimization benchmarks, AlphaOPT steadily improves as more training data become available (65% $\rightarrow$ 72% from 100 to 300 training items) and outperforms the strongest baseline by 9.1% and 8.2% on two out-of-distribution datasets. These results demonstrate that structured experience learning, grounded in solver feedback, provides a practical alternative to retraining for complex reasoning tasks requiring precise formulation and execution. AlphaOPT code and data are available at https://github.com/Minw913/AlphaOPT.
Large Language Models, Optimization Problems, Experience Learning
$c$s: Applied computing Operations researchccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Optimization models support critical decision-making in finance, manufacturing, marketing, transportation, and logistics (AhmadiTeshnizi et al., 2024; Bertsimas and Tsitsiklis, 1997; Ramamonjison et al., 2022). Beyond improving efficiency, automating the optimization workflow lowers the barrier to operations research expertise in industry and enables non-experts to prototype faster, iterate on formulations, and deploy solver-backed decisions at scale. Yet this process has long been challenging, because informal and often ambiguous specifications must be mapped to precise, domain-specific formulations and paired with appropriate code and solvers. This creates major bottlenecks for end-to-end automation (Jiang et al., 2025).
@@
Advances in large language models (LLMs) make this vision increasingly feasible: They can parse natural language requirements (Ouyang et al., 2022), generate executable programs (Nijkamp et al., ; Jimenez et al., 2024), and orchestrate downstream tools (Qin et al., 2024). Two main lines of work have emerged. Prompt-based systems steer general LLMs with structured prompts and tool use (Xiao et al., 2023; Thind et al., 2025; AhmadiTeshnizi et al., 2024; Zhang and Luo, 2025). Fine-tuning approaches adapt models on domain corpora and benchmarks (Huang et al., 2025; Yang et al., 2024). Despite this progress, both families face important limitations. Prompt-based systems improve only within the scope of their fixed prompt templates and are sensitive to small wording changes and domain shifts. Fine-tuned models, on the other hand, require retraining as new data arrives and, critically, most benchmarks and datasets in the community (e.g., NLP4LP (AhmadiTeshnizi et al., 2024), MAMO (Huang et al., 2024), IndustryOR (Huang et al., 2025)) provide only final programs or solutions rather than the intermediate reasoning that governs modeling decisions, which limits the generalizability of fine-tuning approaches.
This motivates a new learning paradigm for optimization formulation: Instead of relying solely on prompts or retraining, LLMs should continually improve by accumulating, refining, and reusing solver-verified modeling insights. Indeed, recent approaches such as Reflexion (Shinn et al., 2023), STaR (Zelikman et al., 2022), ExpeL (Zhao et al., 2024), and AlphaEvolve (Novikov et al., 2025) demonstrate that large models can improve through experiential reuse, storing reflections, rationales, or code edits and applying them in new tasks. These methods have been effective in common reasoning tasks such as information retrieval (Yang et al., 2018) and web browsing (Yao et al., 2022), but they face limitations for optimization problems. First, their experiences are largely preserved as free-form text or edits, without explicit applicability semantics or refinement mechanisms to represent, verify, and refine applicability-sensitive modeling knowledge; in optimization tasks, applying mismatched experiences can therefore have detrimental effects. Second, when generating experiences, these methods often lack rigorous replay-based validation, or rely on verifications limited to task-level outcomes (e.g., rewards, final answers, or test-case success), which does not guarantee that the underlying knowledge is structurally valid or transferable. Third, their experiences tend to be organized as a flat list of items, whereas operations research involves diverse insights across different modeling stages and problem types, inherently requiring a more efficient organization of acquired knowledge.
Regarding these challenges, we introduce AlphaOPT, an experience learning framework for optimization formulation that builds and continually refines a library of solver-verified modeling insights. In contrast to prior approaches that store experience as unstructured text or static edits, AlphaOPT formulates experience as structured, reusable insights with explicit applicability semantics. Each insight is defined as a four-tuple(taxonomy, condition(applicability), explanation, example), specifying not only what modeling rule to reuse, but when and why it should be applied given a natural-language problem description. AlphaOPT improves performance through a continual two-phase cycle without updating model parameters. In the Library Learning phase, the framework self-explores optimization tasks and autonomously constructs a hierarchical taxonomy to organize insights across modeling stages, while relying on solver verification to ensure correctness even in the absence of gold-standard programs. In the Library Evolution phase, AlphaOPT diagnoses systematic misalignments between tasks and insights and refines applicability conditions based on aggregate evidence, preventing insights from being either overly narrow or overly general. Compared to existing experience learning methods, AlphaOPT introduces three key advances: (1) a self-updating hierarchical taxonomy for organizing insights; (2) solver-verified insight generation and updating; and (3) an aggregate-driven refinement mechanism that continuously improves the precision and generalizability of insight applicability.
@@
We conduct quantitative experiments across multiple benchmarks and baselines. Compared with other approaches, the results show that AlphaOPT not only achieves state-of-the-art performance on multiple benchmark datasets without any gold-standard formulation or program, but also demonstrates strong generalization on out-of-distribution datasets, and continual performance growth with converging library size. These results demonstrate the efficacy of experience learning for optimization formulation, which paves the way toward more challenging settings, such as efficient program formulation and large-scale optimization. Compared to existing approaches, AlphaOPT makes the following contributions:
  • We introduce the first experience learning framework for natural-language optimization modeling, which represents solver-verified modeling knowledge as condition-aware insights that are continually refined over time.
@@
  • Our experiments show that AlphaOPT generalizes strongly beyond training distributions, achieving state-of-the-art performance on the LogiOR and OptiBench benchmarks.
  • We provide extensive qualitative and quantitative analyses of the learned experience library and its refinement dynamics, demonstrating explicit, interpretable, and human-inspectable knowledge transfer in operations research domains.
Figure 1. AlphaOPT iteratively learns an experience library through two phases: a Library Learning phase that extracts structured insights from failed attempts and organizes them in a hierarchical taxonomy, and a Library Evolution phase that refines insight applicability based on all associated tasks avoid over- or under-generalization.
The figure shows a two-phase process for learning and refining optimization knowledge. On the left, Phase 1 illustrates library learning. The process begins with an optimization task, including a natural language problem statement, a known correct answer, and a gold-standard mathematical program. The system attempts to solve the task, learns from both mistakes and successful solutions, and uses these attempts to extract reusable experience. This experience is organized into a structured taxonomy that covers domain modeling, general mathematical formulation, and code implementation. From this structure, the system generates concrete insights, each with a clear condition describing when it applies, an explanation of the modeling principle, and examples that contrast incorrect and correct formulations. On the right, Phase 2 illustrates library evolution. Previously generated insights are examined by comparing cases where they apply and where they do not. Using performance and error signals aggregated across tasks, the system refines the conditions under which each insight should be used. The result is a more precise description of applicability, for example distinguishing problems that require closed vehicle tours with flow balance from problems that allow open or multi-day routes. Overall, the figure shows how raw optimization attempts are transformed into structured, conditional knowledge and how that knowledge is iteratively refined to improve preciseness of the applicability and transfer across diverse problems.
2. Related Work
2.1. LLMs for Solving Optimization Problems
Recent work increasingly uses LLMs as autonomous agents for optimization, mainly following two paradigms: formulation and heuristic generation (Ye et al., 2024; Yang et al., 2025b; Wu et al., 2025). LLM-driven formulation tasks aim to translate natural-language problem statements into solver-ready mathematical programs and executable solving code. Multiple benchmark datasets have been proposed to cover diverse problem families such as LP, MILP, and NLP (Xiao et al., 2023; AhmadiTeshnizi et al., 2024; Yang et al., 2024, 2025a).
Existing approaches can be broadly grouped into prompt-based and training-based methods. Prompt-based methods use multi-step prompting and structured reasoning to guide LLMs in parsing problems, extracting variables/constraints, and synthesizing models. Representative examples include Chain-of-Experts (Xiao et al., 2023), which leverages multi-role decomposition, and OptiMUS (AhmadiTeshnizi et al., 2024), which employs a self-reflect agentic pipeline with retrieval-augmented prompting to produce mathematical formulations. LEAN-LLM-OPT (Liang et al., 2026) further integrates a three-agent workflow with RAG and few-shot learning for problem typing, example retrieval, and data grounding. However, these methods lack sustained knowledge injection and long-term memory, which can lead to performance plateaus under distribution shift. Training-based methods improve robustness by building instruction corpora and fine-tuning open-source models. For example, ORLM (Huang et al., 2025) synthesizes natural-language optimization problem data used for fine-tuning, and LLMOPT (Jiang et al., 2024) adopts a five-element schema with multi-instruction SFT on expert-labeled and augmented data. Recent post-training work further applies RL-style refinement, such as OR-R1 (Ding et al., 2025), to improve single-shot performance and robustness. Nonetheless, these pipelines rely heavily on high-quality, diverse data, and are often hard to interpret or audit.
Overall, despite progress in prompting, data synthesis, fine-tuning, and post-training, existing methods remain insufficient for reliable deployment and sustained improvement in real-world operations research systems.
2.2. Decision-making Tasks with Experience Learning
Experience Learning studies how LLM-based agents extract and reuse transferable structured experience from historical interactions to continuously improve downstream decision-making and execution (Zhao et al., 2023; Shinn et al., 2023). Such experience can take the form of episodic traces (Park et al., 2023) (e.g., tool/action sequences), semantic insights abstracted from successful or failed attempts (Zhao et al., 2023). This line of work is primarily evaluated on long-horizon, interactive, feedback-rich decision-making tasks, including multi-tool planning and web interaction (Qin et al., 2023; Zhou et al., 2023), repository-level program repair/code generation (Jimenez et al., 2023), and multi-step reasoning problem solving (Shinn et al., 2023; Wang et al., 2024).
Shinn et al. (2023) propose Reflexion, which enables language agents to learn from trial-and-error by converting task feedback into verbal self-reflections that are stored as episodic memory and reused in subsequent attempts. ExpeL (Zhao et al., 2023) jointly extracts similar exemplars/trajectories together with their associated insights to guide task execution and further employs a voting-based mechanism to downweight mutually contradictory insights, thereby improving single-shot performance and cross-task generalization. At a more symbolic end, Zhu et al. (2023) explicitly induces a rule library and maintains it using quantitative filters such as coverage and confidence, enhancing controllability and robustness of experience reuse. AgentKB (Tang et al., 2025) and ReasoningBank (Ouyang et al., 2025) move toward organizing an evolving reasoning memory to support self-improving agents across tasks.
However, current frameworks face systematic limitations in three key aspects. First, regarding experience applicability, most methods still rely on embedding-based similarity retrieval or coarse metadata to match experiences (Zhao et al., 2023; Park et al., 2023; Wang et al., 2024), while lacking explicit, verifiable preconditions (when an experience applies, when it should not, and what assumptions it requires), which increases the risk of negative transfer under distribution shift. Second, for strict verification of experience generation, merging, and editing, existing mechanisms often depend on LLM self-assessment (Zhao et al., 2023; Shinn et al., 2023; Feng et al., 2025); for instance, ExpeL’s voting-based filtering can remove contradictions but does not guarantee semantic correctness. Finally, regarding cross-task generalization, existing experience libraries typically support experience accumulation and local pruning (Zhao et al., 2023; Tang et al., 2025; Ouyang et al., 2025), but lack a systematic mechanism to assess whether an experience that helps on local source tasks may degrade performance on broader related task sets.
3. Methodology
@@
Building reliable LLM-based systems for solving optimization problems remains challenging. To begin with, gold-standard formulations and programs are scarce and may contain annotation errors (Jiang et al., 2025; Yang et al., 2025a), while datasets with only answer-level supervision are underexploited (Huang et al., 2024, 2025; Lu et al., 2025). Moreover, although optimization tasks arrive with diverse natural-language descriptions, they often share recurring modeling rules that activate under identifiable conditions; however, existing methods still struggle to learn such when-to-apply-what knowledge. Fine-tuned models (Huang et al., 2025; Jiang et al., 2025) often overfit to syntax without mastering applicability, whereas prompt-based agent systems (AhmadiTeshnizi et al., 2024; Xiao et al., 2023; Yang et al., 2025a) rely on human empirical curation and lack the capacity to adapt or continually learn from larger datasets.
To this end, we propose AlphaOPT, which constructs an experience library of reusable insights with explicit applicability conditions, and then iteratively evolves these conditions at population level to improve generalization. As illustrated in Figure 1, it consists of two complementary phases that form a continual cycle of acquisition and refinement. The first phase, Library Learning, extracts structured insights from individual failed tasks under answer-only supervision while minimizing redundancy. The second phase, Library Evolution, diagnoses misalignments between insights and tasks and refines applicability conditions to enhance generalization while reducing confusion caused by overgeneralization. In Section 3.3, we provide a mathematical interpretation that frames library construction as maximizing task success with a size regularizer. Implementation details for the different phases are provided in the Appendix A.
To the best of our knowledge, our framework is the first to adapt experience learning to operations research (OR) with three key innovations: (1) structured knowledge for interpretability and auditability: Each insight is represented with taxonomy, condition, explanation, and example, which renders its applicability explicit, reviewable, and even revisable in practice; (2) solver-guided verifiability: If a program generated by LLM achieves the optimal objective under the solver, it is highly likely to be valid and can serve as a reliable anchor for extracting insights, which broadens the sources of experience collection. Newly generated insights are retest to be retrievable in the library and to succeed when replayed on its source task, which ensures that they are valid before integration; (3) refinement of experience applicability for generalizability and preciseness: Applicability conditions are refined using cross-task evidence, so insights neither over-generalize nor remain too narrow, which improves safe transfer across problem families.
@@
Figure 2. The overall workflow of AlphaOPT. The left panel depicts two complementary flows: library learning, which self-explores correct patterns and extracts new insights by trial and error; library diagnosis, which categorizes interactions between failed tasks and retrieved insights into positive, negative, and unretrieved cases. The right panel shows library refinement, whereby the LLM refines insight applicability, validates via retrieval replay, and reintegrates them into the library.
The figure illustrates how the insight library is learned, diagnosed, and refined through interaction with optimization tasks. On the left, the library learning and diagnosis process is shown. Given an optimization task, the system retrieves relevant prior insights based on a taxonomy and their applicability conditions. These insights guide the construction of a mathematical formulation, which is then translated into a solving program. If the program succeeds, the task provides positive evidence for the retrieved insights. If it fails, the system performs self-exploration by comparing the failed solution with a gold-standard program and extracting new insights. All insights, both existing and newly generated, are stored in the library. At the same time, the system diagnoses insight applicability across related tasks. For each task, an insight can be observed as helpful, harmful, or not retrieved at all. Insights that consistently perform poorly or are missing when needed are marked as candidates for refinement. On the right, the library refinement process is shown. Candidate insights are replayed through library retrieval and self-verification on source tasks. Based on positive, negative, or unretrieved outcomes, their applicability conditions are updated by keeping, removing, or adding conditions. The refined versions are validated, and the best-performing versions are selected to form the updated library. Overall, the figure shows a closed loop in which solving optimization tasks both uses the library and continuously improves it by refining when each insight should apply.
3.1. Library Learning